# TFLAG: Towards Practical APT Detection via Deviation-Aware Learning on Temporal Provenance Graph

### v1

@@ VI EXPERIMENTS AND ANALYSIS
Our system was developed using Python. Log parsers were constructed for various datasets, and a unified format for graph access was implemented in the database. The graph learning framework utilizes a multi-layer TGAT [23] and PyTorch [39]. Evaluations were conducted across four public datasets, employing two detection strategies: graph-level and window-level. All experiments were performed on a server running Ubuntu 22.04.4 LTS with kernel version 5.15.0-117-generic, equipped with an Intel Xeon Platinum 8352V processor (16 cores at 2.10 GHz) and 64 GB of RAM. Within the TFLAG system settings, the graph encoder’s output dimension is set at 128, the training batch size at 512, the number of neighbor nodes at 20, and the learning rate at 0.0001. Subsequent experiments examined the influence of varying hyperparameters on performance, with selections including N from 5, 10, 20, 30, 40, h from 32, 64, 128, 256, e from 4, 8, 16, 32, and T from 5, 10, 15, 20. In Section V-A, we introduce the datasets utilized by the TFLAG system at both graph and window levels. Section V-B details the system’s effectiveness across different scenarios, comparing it to the latest research on the same granularity. Section V-C provides additional comparisons, while Section V-D offers an in-depth comparison of the impact of various TFLAG hyperparameters on the system’s detection outcomes. The entire set of experiments robustly demonstrates the TFLAG system’s precision in detecting anomalies under unsupervised conditions.
@@ VI-A Datasets
We evaluate the effectiveness of TFLAG on multiple public datasets: StreamSpot [40], CADETS, CLEARSCOPE, and THEIA datasets in DARPA Engagement3 [41]. These datasets vary in granularity, logging formats, attack methods, and sources. Verifying the effectiveness of the system on these diverse datasets can reflect the performance of TFLAG, and currently the most advanced APT detection methods are also tested on these datasets, allowing for effective result comparisons. The composition and source format of the datasets are introduced in detail below.
StreamSpot Datasets. This dataset provides provenance graph data by simulating a laboratory environment, capturing normal and attack behaviors across six scenario graphs. Each scenario graph comprises 100 graphs that record the underlying process traffic monitored in a time series. As detailed in Table I, five of these scenarios (YouTube, Gmail, Video Game, Attack, Download, CNN) represent normal user behaviors and are entirely benign; the remaining scenario consists of 100 graphs depicting abnormal behavior generated by simulating driver download attacks. It is important to note that this dataset focuses on graph-level anomaly detection rather than fine-grained window or entity-level detection. Following its previous methodologies, the TFLAG method performs graph-level detection on this dataset. Each scenario graph is composed of chronologically arranged edges, aligning with the TFLAG premise of training on benign datasets and testing on abnormal ones. For our approach, one graph from each benign scenario is used for training, while the rest of the benign graphs and all abnormal graphs are utilized for testing and evaluation. This method enables us to determine whether each graph is anomalous and to benchmark our detection results against current SOTA APT detection methods.
TABLE I: Overvier of DARPA-E3 Dataset
@@
CADETS Nginx Backdoor 613,713 29,727,532 17.9
THEIA Browser Extension 1,259,368 44,066,112 81.9
ClearScope Firefox Backdoor 201,021 8,915,438 16.6
DARPA E3 Datasets. The DARPA dataset originates from the third Transparent Computing (TC) adversarial exercise. It encompasses dataset generation techniques, scenario descriptions, tools used, attack details, evaluations of each analysis report, sample attack graphs, and logs. The attacker’s goal was to steal proprietary and personal information from a target company by exploiting a FreeBSD web server, injecting the SSHD process, and monitoring network activity to exploit discovered hosts and extract data. Ultimately, the attacker targeted a mobile device to extract personally identifiable information (PII), using tools such as Nginx and Firefox backdoors, browser extensions, and Drakon APT. Additionally, the attack involved HTTP communication, creating elevated processes, and injecting .dll or .so files. Multiple logging systems captured host activity, documenting underlying process traffic throughout the exercise.
TABLE II: Dataset Label Division
@@
CADETS 2018-04-03/04/05 2018-04-07 2018-04-06
THEIA 2018-04-03/04/05 2018-04-09 2018-04-10
ClearScope 2018-04-07/08/09 2018-04-10 2018-04-11
We evaluated the TFLAG method using the CADETS, CLEARSCOPE, and THEIA sub-datasets [41], extracting critical edges and nodes essential for model training and anomaly detection. The combined size of these datasets is 124.2GB, as shown in Table II. To assess the model’s entity-level detection capabilities on this large, unlabeled dataset effectively, we implemented a daily training regimen. Specifically, with the CADETS dataset, we preprocessed and trained using the first two days of normal data. We used the third day’s benign data for evaluation and the fourth day’s data, which included abnormal attacks, for testing. This approach allowed us to evaluate authentic attack behaviors and benchmark detected abnormal event windows against the E3 dataset’s ground truth, enabling us to accurately identify the critical exceptional window and correlate all related anomalous events without prior knowledge.
As highlighted in Fig. 5, the E3 dataset provides detailed information on attack events, including specific timestamps and behavioral descriptions. This data facilitates the precise alignment of detected anomalous event windows with actual attack periods. Unlike previous methods that relied on manual annotation [42], our approach does not require predefined label information and can independently identify the core abnormal activity window solely by learning from benign data. Additionally, by performing anomaly detection on the DARPA dataset, we can compare our results with the current SOTA systems in anomaly-based provenance graph detection, thus validating the effectiveness of the TFLAG method.
@@
TABLE III: RESULTS OF WINDOW LEVEL DETECTION EXPERIMENTS IN DARPA-E3 DATASETS
@@
CADETS 4 181 0 0 1.000 1.000 1.000 1.000
ClearScope 4 180 0 1 1.000 0.800 0.995 0.900
THEIA 7 117 0 1 1.000 0.875 0.992 0.938
TABLE IV: TFLAG and KAIROS examine the distinctive attributes of the most anomalous edges within the same anomaly window and compare them with actual attack scenarios.
@@
CADETS F1-elevate /tmp/vUgefa KAIROS 4-6-11:33$\sim$4-6-11:48 netflow:128.55.12.10:53 → subject: lsof 7.4449
CADETS F1-elevate /tmp/vUgefa TFLAG 4-6-11:30$\sim$4-6-11:45 subject:vUgefal → file:/tmp/vUgefal 23.4677
VI-B Effectiveness
We evaluated the effectiveness of the TFLAG method for detecting APT attacks across four datasets at varying levels of granularity. To emulate real-world scenarios, training was conducted on days without anomalies in chronological order for each dataset, using the days with anomalies and their adjacent days as the subjects of detection. Detailed results of these analyses are presented in Table III , where they are compared with the performance metrics of leading APT detection methods under the same conditions.
Precision, recall, accuracy, and the area under the receiver operating characteristic curve (AUC) for the datasets were computed based on predefined time windows. The attack log files provided were manually annotated to differentiate between bengin and attack-related time windows. If a time window was inaccurately labeled as anomalous by the TFLAG method, it was classified as a false positive (FP); conversely, correct identification of an attack window by the TFLAG method was considered a true positive (TP). The instances of false negatives (FN) and true negatives (TN) were calculated using the same rationale. Table III also enumerates the specific counts of TP, TN, FP, and FN for each dataset. The results indicate that the TFLAG method successfully identified all abnormal windows but exhibited false negatives in the ClearScope and THEIA datasets, the details of which are further dissected below.
Upon further examination of the missed time windows, we identified specific limitations in the TFLAG method’s detection capabilities within the THEIA and ClearScope datasets. In THEIA, the TFLAG method erroneously labeled the initial time window of the attack sequence as benign. A detailed log analysis revealed that this misclassification occurred as the attacker attempted to deploy attack code at the www.allstate.com website to deliver a malicious payload. A system lock-up caused the system to become unresponsive, delaying detection and preventing access via SSH or VNC. During this period, the system continued to emit stale data, leading TFLAG to incorrectly classify this timeframe as benign. Once the system regained stability and resumed normal operations, subsequent attack efforts were correctly logged.
In ClearScope, TFLAG failed to recognize the final time window of the attack as malicious. During this window, the attacker made unsuccessful attempts to perform privileged file operations, which ultimately led to the termination of the remote attack. The system then returned to normal operation for the remainder of the window without any further malicious activity. The dominance of normal activity throughout this period led TFLAG to mistakenly classify it as non-anomalous. These detection errors by TFLAG can generally be attributed to its sensitivity to the timing and extent of manifest attack behaviors within the designated windows.
(a) $N$
@@ VI-B Effectiveness
VI-C TFLAG vs. SOTA.
The datasets used to evaluate the TFLAG method are similarly applied to assess SOTA APT detection methods. Both TFLAG and KAIROS employ a label-free, self-supervised anomaly detection system, as opposed to the label-based PIDS[6, 43, 44, 45]. Therefore, direct comparisons with more advanced methodologies such as FLASH, MAGIC, HOLMES, and JBEIL, which depend on project code analysis and abnormal node labels from TreaTrace or manually generated labels, are impractical. These labels significantly affect detection quality. Although numerous studies focus on anomaly-based APT detection, many are proprietary and difficult to replicate. According to its publications[13], KAIROS has surpassed these methods in terms of detection efficiency. Consequently, this study concentrates on comparing TFLAG with KAIROS to ascertain the effectiveness of the TFLAG method in detecting APT attacks. This comparison clarifies TFLAG’s performance and advantages in unsupervised APT detection, independent of labeled dataset quality control concerns.
KAIROS utilizes an encoder-decoder framework within a Temporal Graph Neural Network (TGN) to detect anomalies by reconstructing network edges, encoding neighborhood structures, and updating them via a memory module. Reproduction of the code revealed that, during its detection phase, numerous false positives associated with system behavior result in a higher recall rate but a reduced accuracy rate. In contrast, TFLAG shows higher precision in identifying anomalous edges, closely aligning with the anomaly events documented in the E3 attack dataset, as detailed in Table IV. Under identical training conditions and window size, TFLAG precisely identifies the malicious subject vUgefal writing to the file /tmp/vUgefal within the attack window, recording a peak reconstruction loss of 23.5. This clearly differentiates it from other events in the same window. In contrast, KAIROS erroneously classifies benign system behavior—specifically, the lsof process checking the network 128.55.12.10:53, which the attack report does not mention—as an attack event, due to its reconstruction loss being the highest in the entire window at 7.4. Additionally, TFLAG can discern incidental events within the window by considering the deviation loss of anomaly scores, rather than solely relying on reconstruction loss for classification. This distinction is substantial because, while KAIROS can detect APT behaviors by assessing the overall loss across multiple windows, its tendency to misclassify normal activities as anomalies results in failing to detect the subtle yet continuous perturbations characteristic of covert attacks. In contrast, TFLAG consistently and accurately identifies genuine attacks, reduces false positives, and thus delivers more reliable anomaly detection outcomes.
Table V demonstrates that TFLAG consistently surpasses KAIROS in accuracy across all three window detection levels. Moreover, in the node detection level dataset SteamSpot, both TFLAG and KAIROS achieve an exceptional 100% accuracy and recall. In the E3 dataset, TFLAG correctly identified the entire attack event within the CADETS anomaly detection window, achieving 100% accuracy. Conversely, KAIROS erroneously categorized a prior window as anomalous, detracting from its accuracy. Similar observations were made in the ClearScope and THEIA datasets; although TFLAG occasionally misses anomaly windows at the beginning or end of an attack, it does not erroneously identify benign windows as anomalous, primarily due to its enhanced sensitivity to anomaly edges compared to KAIROS. Overall, TFLAG’s methodology for identifying anomaly windows far surpasses that of KAIROS, reducing false positives and improving both recall and accuracy rates, thus proving more effective and robust in APT detection.
TABLE V: Comparison study between TFLAG and KAIROS

### v2

@@ VI EXPERIMENTS AND ANALYSIS
Our system was developed using Python. Log parsers were constructed for various datasets, and a unified format for graph access was implemented in the database. The graph learning framework utilizes a multi-layer TGAT [23] and PyTorch [38]. Evaluations were conducted across four public datasets, employing two detection strategies: graph-level and window-level. All experiments were performed on a server running Ubuntu 22.04.4 LTS with kernel version 5.15.0-117-generic, equipped with an Intel Xeon Platinum 8352V processor (16 cores at 2.10 GHz) and 64 GB of RAM. Within the TFLAG system settings, the graph encoder’s output dimension is set at 128, the training batch size at 512, the number of neighbor nodes at 20, and the learning rate at 0.0001. Subsequent experiments examined the influence of varying hyperparameters on performance, with selections including neighbor size N from 5, 10, 20, 30, 40, hidden layer dimension h from 32, 64, 128, 256, edge embedding dimension e from 4, 8, 16, 32, and time window size T from 5, 10, 15, 20. In Section V-A, we introduce the datasets utilized by the TFLAG system at both graph and window levels. Section V-B details the system’s effectiveness across different scenarios, comparing it to the latest research on the same granularity. Section V-C provides additional comparisons, while Section V-D offers an in-depth comparison of the impact of various TFLAG hyperparameters on the system’s detection outcomes. The entire set of experiments robustly demonstrates the TFLAG system’s precision in detecting anomalies under unsupervised conditions.
@@ VI-A Datasets
We evaluate the effectiveness of TFLAG on multiple public datasets: StreamSpot [39] and DARPA [40, 41]. These datasets vary in granularity, logging formats, attack methods, and sources. Verifying the effectiveness of the system on these diverse datasets can reflect the performance of TFLAG, and currently the most advanced APT detection methods are also tested on these datasets, allowing for effective result comparisons. The composition and source format of the datasets are introduced in detail below.
TABLE I: Overvier of StreamSpot Dataset
Scenario ALL #Nodes ALL #Edges Size(GB)
Dwnload 8,830 310,814 1.0
Gmail 6,826 37,382 0.1
VGame 8,636 112,958 0.4
YouTube 8,292 113,229 0.3
Attack 8,890 28,423 0.1
CNN 8,989 294,903 0.9
StreamSpot Datasets. This dataset provides provenance graph data by simulating a laboratory environment, capturing normal and attack behaviors across six scenario graphs. Each scenario graph comprises 100 graphs that record the underlying process traffic monitored in a time series. As detailed in Table I, five of these scenarios (YouTube, Gmail, Video Game, Attack, Download, CNN) represent normal user behaviors and are entirely benign; the remaining scenario consists of 100 graphs depicting abnormal behavior generated by simulating driver download attacks. It is important to note that this dataset focuses on graph-level anomaly detection rather than fine-grained window or entity-level detection. Following its previous methodologies, the TFLAG method performs graph-level detection on this dataset. Each scenario graph is composed of chronologically arranged edges, aligning with the TFLAG premise of training on benign datasets and testing on abnormal ones. For our approach, one graph from each benign scenario is used for training, while the rest of the benign graphs and all abnormal graphs are utilized for testing and evaluation. This method enables us to determine whether each graph is anomalous and to benchmark our detection results against current state-of-the-art APT detection methods.
TABLE II: Overvier of DARPA Dataset
@@
CADETS-E3 Nginx Backdoor 613,713 29,727,532 17.9
THEIA-E3 Browser Extension 1,259,368 44,066,112 81.9
ClearScope-E3 Firefox Backdoor 201,021 8,915,438 16.6
CADETS-E5 Nginx Drakon 394,617 77,886,133 106.9
THEIA-E5 Firefox Drakon 1,511,241 74,565,731 140.3
ClearScope-E5 Lockwatch APK 268,251 34,413,590 67.7
DARPA Datasets. The DARPA dataset originates from the third and fifth Transparent Computing (TC) adversarial exercises, as shown in Table II. It encompasses dataset generation techniques, scenario descriptions, tools used, attack details, evaluations of each analysis report, sample attack graphs, and logs. The attacker’s goal was to steal proprietary and personal information from a target company by exploiting a FreeBSD web server, injecting the SSHD process, and monitoring network activity to exploit discovered hosts and extract data. Ultimately, the attacker targeted a mobile device to extract personally identifiable information (PII), using tools such as Nginx and Firefox backdoors, browser extensions, and Drakon APT. Additionally, the attack involved HTTP communication, creating elevated processes, and injecting .dll or .so files. Multiple logging systems captured host activity, documenting underlying process traffic throughout the exercise.
TABLE III: Dataset Label Division
@@
CADETS-E3 2018-04-03/04/05 2018-04-07 2018-04-06
THEIA-E3 2018-04-03/04/05 2018-04-09 2018-04-10
ClearScope-E3 2018-04-07/08/09 2018-04-10 2018-04-11
CADETS-E5 2019-05-08/09/11 2019-05-17 2019-05-16
THEIA-E5 2019-05-08/09/11 2019-05-14 2019-05-15
ClearScope-E5 2019-05-08/09/11 2019-05-15 2019-05-17
We evaluated the TFLAG method using the underlying log data obtained from different capture systems [40] in the adversarial exercises, extracting critical edges and nodes essential for model training and anomaly detection. As shown in Table III, to assess the model’s entity-level detection capabilities on this large, unlabeled dataset effectively, we implemented a daily training regimen. Specifically, with the CADETS-E3 dataset, we preprocessed and trained using the first two days of normal data. We used the third day’s benign data for evaluation and the fourth day’s data, which included abnormal attacks, for testing. This approach allowed us to evaluate authentic attack behaviors and benchmark detected abnormal event windows against the DARPA dataset’s ground truth, enabling us to accurately identify the critical exceptional window and correlate all related anomalous events without prior knowledge.
As highlighted in Fig. 5, the DARPA dataset provides detailed information on attack events, including specific timestamps and behavioral descriptions. This data facilitates the precise alignment of detected anomalous event windows with actual attack periods. Unlike previous methods that relied on manual annotation [42], our approach does not require predefined label information and can independently identify the core abnormal activity window solely by learning from benign data. Additionally, by performing anomaly detection on the dataset, we can compare our results with the current SOTA systems in anomaly-based provenance graph detection, thus validating the effectiveness of the TFLAG method.
@@
VI-B Effectiveness
We evaluated the effectiveness of the TFLAG method for detecting APT attacks across multiple scenario datasets at varying levels of granularity. To emulate real-world scenarios, training was conducted on days without anomalies in chronological order for each dataset, using the days with anomalies and their adjacent days as the subjects of detection. We conducted experiments on the SteamSpot dataset and DARPA, and the detailed results of these analyses are presented in Table IV, where the evaluation metrics used are consistent with those commonly employed by other methods in the field.
Precision, recall, accuracy, and the area under the receiver operating characteristic curve (AUC) for the datasets were computed based on predefined time windows. The attack log files provided were manually annotated to differentiate between bengin and attack-related time windows. If a time window was inaccurately labeled as anomalous by the TFLAG method, it was classified as a false positive (FP); conversely, correct identification of an attack window by the TFLAG method was considered a true positive (TP). The instances of false negatives (FN) and true negatives (TN) were calculated using the same rationale. Table IV also enumerates the specific counts of TP, TN, FP, and FN for each dataset. The results indicate that the TFLAG method successfully identified all abnormal windows in the DARPA-E3 and StreamSpot datasets, but exhibited false negatives in the ClearScope-E3 and THEIA-E3 datasets, and false positives in the DARPA-E5 dataset. The details of these are further dissected below.
Upon further examination of the missed time windows, we identified specific limitations in the TFLAG method’s detection capabilities within the THEIA-E3 and ClearScope-E3 datasets. In THEIA-E3, TFLAG erroneously labeled the initial time window of the attack sequence as benign due to a system lock-up, which caused the system to become unresponsive and delayed detection. During this period, the attacker attempted to deploy attack code at the www.allstate.com website to deliver a malicious payload, but the system continued to emit stale data, leading TFLAG to incorrectly classify the timeframe as benign. Once the system regained stability and resumed normal operations, subsequent attack efforts were correctly logged. Similarly, in ClearScope-E3, TFLAG failed to recognize the final time window of the attack, during which the attacker made unsuccessful attempts at privileged file operations. Although the system returned to normal after the attack, the dominance of normal activity in this window led TFLAG to mistakenly classify it as non-anomalous. These detection errors by TFLAG can generally be attributed to its sensitivity to the timing and extent of manifest attack behaviors within the designated windows. However, it is important to note that these issues do not impact the administrator’s ability to detect the attack or the overall APT attack analysis, as the attack sequence was eventually detected, and the overall analysis remains intact.
TABLE IV: Detection Experiment Results of Selected Datasets
@@ VI-B Effectiveness
CADETS-E3 4 181 0 0 1.000 1.000 1.000 1.000
ClearScope-E3 4 180 0 1 1.000 0.800 0.995 0.900
THEIA-E3 7 117 0 1 1.000 0.875 0.992 0.938
CADETS-E5 6 246 3 1 0.667 0.857 0.984 0.923
ClearScope-E5 10 240 4 1 0.714 0.909 0.800 0.946
THEIA-E5 2 180 1 0 0.667 1.000 0.994 0.997
TABLE V: TFLAG and KAIROS examine the distinctive attributes of the most anomalous edges within the same anomaly window and compare them with actual attack scenarios.
@@
CADETS F1-elevate /tmp/vUgefa KAIROS $3\sim4-6-11$:48 netflow:128.55.12.10:53 → subject: lsof 7.4449
[…]
In the DARPA-E5 scenario, the attacker initially exploited a vulnerable process before proceeding to target additional processes. Consequently, previously attacked components were mistakenly identified as attack events upon resuming normal system behavior. As an anomaly-based detection system, TFLAG demonstrates improved sensitivity to both false positives and genuine attacks that deviate from established norms. This resulted in a misclassification, wherein previously compromised events were erroneously categorized as normal behavior. To effectively mitigate false positives associated with normal system behavior, it is essential to implement re-labeling and retraining in subsequent stages.
(a) Kairos reconstruction loss distribution
(b) TFLAG reconstruction loss distribution
Figure 6: The distribution of the highest reconstruction loss events for each time window in the detection results of 2018-04-06, for both Kairos and TFLAG in CADETS-E3 is presented.
Figure 7: Comparison of true anomalous events among the top 10% reconstruction loss events for TFLAG and KAIROS within the same anomalous time window.
VI-C TFLAG vs. SOTA.
The datasets used to evaluate the TFLAG method are similarly applied to assess SOTA APT detection methods. Both TFLAG and KAIROS employ a label-free, self-supervised anomaly detection system, as opposed to the label-based PIDS[6, 43, 44, 45]. Therefore, direct comparisons with more advanced methodologies such as FLASH, MAGIC, HOLMES, and JBEIL, which depend on project code analysis and abnormal node labels from TreaTrace or manually generated labels, are impractical. While numerous studies examine anomaly-based detection of APT, the detection granularity and labeling used for comparison in their systems differ from those in our approach. According to its publications[13], KAIROS has surpassed these methods in terms of detection efficiency. Consequently, this study concentrates on comparing TFLAG with KAIROS to ascertain the effectiveness of our method in detecting APT attacks. This comparison clarifies TFLAG’s performance and advantages in unsupervised APT detection, independent of labeled dataset quality control concerns.
KAIROS utilizes an encoder-decoder framework within a Temporal Graph Neural Network (TGN) to detect anomalies by reconstructing network edges, encoding neighborhood structures, and updating them via a memory module. Reproduction of the code revealed that, during its detection phase, numerous false positives associated with system behavior result in a higher recall rate but a reduced accuracy rate. In contrast, TFLAG demonstrates higher precision in detecting anomalous edge events. As shown in Fig. 6, within the CADETS-E3 scenario, TFLAG successfully detected the most significant anomalous loss during the attack event window. Specifically, within the first fifteen minutes after the attack began, KAIROS generated a false positive with notable anomalous activity, which was not an actual attack. In the following four time windows, the detected loss values did not show a significant increase. KAIROS later identified the attack queue after filtering out noise nodes and analyzing node rarity, but this was done with the false positive window still present. In contrast, TFLAG accurately captured the high-anomaly events within the attack window, where the anomaly score was significantly higher than the maximum values in other time windows. Additionally, as observed in the figure, TFLAG also successfully identified high-loss events in the previous time window. By analyzing the deviation scores, the rarity of the anomalous events, and the loss values, it effectively filtered out this false-positive time window. This precise anomaly detection method allows TFLAG to provide more reliable attack identification. Table V shows that the anomalous events detected by TFLAG align highly with the actual records in the E3 attack dataset.
Under the same training conditions and time windows, TFLAG accurately identified the malicious entity vUgefal writing the file /tmp/vUgefal within the attack window, recording a peak reconstruction loss of 20.7. This peak clearly distinguished it from other events within the same time window. In contrast, KAIROS erroneously classified a benign system behavior, specifically the lsof process checking the network 128.55.12.10:53, as the most anomalous event. This event was not mentioned in the attack report, but its reconstruction loss was the highest in the entire time window, reaching 7.4. This impacted subsequent administrator trace-back decisions. Furthermore, the construction of the deviation network effectively differentiates between attack events and false alarm events, as illustrated in Fig. 7. Within the same attack window, we can identify a greater number of genuine attack events among the top 10% of anomalous events. This capability allows administrators to more accurately pinpoint rare attack events exhibiting a higher degree of anomaly during subsequent fine-grained window-based trace analysis.
Table VI demonstrates that TFLAG consistently surpasses KAIROS in accuracy across all three window detection levels. Moreover, in the node detection level dataset SteamSpot, both TFLAG and KAIROS achieve an exceptional 100% accuracy and recall. In the E3 dataset, TFLAG correctly identified the entire attack event within the CADETS anomaly detection window, achieving 100% accuracy. Conversely, KAIROS erroneously categorized a prior window as anomalous, detracting from its accuracy. Similar observations were made in the ClearScope and THEIA datasets; although our method occasionally misses anomaly windows at the beginning or end of an attack, it does not erroneously identify benign windows as anomalous, primarily due to its enhanced sensitivity to anomaly edges compared to KAIROS. Overall, TFLAG’s methodology for identifying anomaly windows far surpasses that of KAIROS, reducing false positives and improving both recall and accuracy rates, thus proving more effective and robust in APT detection.
(a) NN
@@
TABLE VI: Comparison study between TFLAG and KAIROS
@@
Figure 9: The distribution of the highest reconstruction loss events for each time window in the detection results of TFLAG under the no deviation network condition in CADETS-E3 of 2018-04-06.